OpenStudy (anonymous):

please help find the vertex and the equation of the axis of symmetry. graph the function f(x)=(x+7)^2

OpenStudy (campbell_st):

ok... so the vertex form of a parabola is \[y = (x - h)^2 + k\] which will have a vertex at (h, k) in your question k = 0 you just need to find h...

OpenStudy (campbell_st):

the axis of symmetry will be x = h

OpenStudy (anonymous):

i dont understand how to find either :(

OpenStudy (anonymous):

y=(x+7)^2 standard form is y=(x-a)^2+b where a and b are the coordinates for vortex. the axis of symmetry is alwasy the linear variable (for parabola) here linear variable is y .